OpenStudy (anonymous):

State what the length of the other leg is in a right triangle whose hypotenuse is 13 and one leg is 12 by using the Pythagorean Theorem.

OpenStudy (anonymous):

6?

OpenStudy (kinggeorge):

6 Is not correct. The pythagorean theorem says that \[c=\sqrt{a^2+b^2}\]correct?

OpenStudy (anonymous):

Yes. I tried that and it doesn't come up with any of the multiple choice answers.

OpenStudy (kinggeorge):

Let's rearrange it to find \(a\).\[\begin{aligned} c&=\sqrt{a^2+b^2}\\ c^2&=a^2+b^2\\ c^2-b^2&=a^2\\ \sqrt{c^2-b^2}&=a \end{aligned}\]

OpenStudy (kinggeorge):

Now plug in \(13=c\), and \(12=b\). We get \[\sqrt{13^2-12^2}=\sqrt{169-144}=\sqrt{25}\]So what's \(\sqrt{25}\)?

OpenStudy (anonymous):

Is it simplified to 5? And the other leg.

OpenStudy (kinggeorge):

Right. So the length of your third leg must be 5.

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (kinggeorge):

No problem.